OpenStudy (anonymous):

Essay: Show all work. A gardener wants to create a triangular garden in the shape of a right triangle with shortest side length x – 5y ft. and the middle length side x + 6y ft. What is an algebraic expression for the area of the garden? Be sure to multiply this out and express in simplest correct mathematical form, including units.

OpenStudy (saifoo.khan):

you have to write that for 50 words? Right?

OpenStudy (chaise):



OpenStudy (chaise):

Is this right?

OpenStudy (chaise):

Please confirm if this is right, so I can help you further?

OpenStudy (anonymous):

The hypotenuse is always the longest side. Your picture is correct chaise.

OpenStudy (chaise):

It says "the shortest side is: x-5y" and the middle length is: x+6y. Is the middle length the hypotenuse?

OpenStudy (anonymous):

I'm not sure if this is right or not. I'm seeking help

OpenStudy (chaise):

I was hopping you had a diagram jbmims.

OpenStudy (anonymous):

Wish I did. Would you multiply each side?

OpenStudy (anonymous):

Think a^2 + b^2 = c^2. The two equations describe a and b (which are the two shorter sides). The "shorter side" corresponds to a and the "middle side" corresponds to b. C is the hypotenuse, so you're correct.

OpenStudy (anonymous):

The shortest side = x - 5*y = height of triangle The middle side = x + 6 y = base of the triangle The longest side = hypotenuse Substituting the above values into the equation for the area of a triangle Area of a triangle = base * height/2 Area of the garden = (x + 6*y)(x - 5y)/2 in feet^2